A laser printer uses a laser with a wavelength of 805nm. What should be the focal length of the lens to focus a collimated laser beam of a diameter of 5mm to obtain a printing resolution of 600 DPI (dots per inch)?


1
Expert's answer
2021-04-30T11:23:30-0400

"\Gamma=\frac Dd=\frac{\lambda}{F},"

"F=\frac{d\lambda}{D},"

"F=\frac{5\cdot 805}{43}=9.4~cm."